OpenStudy (anonymous):

help i need somebody to chech my questions im not 100% sure i got them all right

OpenStudy (anonymous):

check*

OpenStudy (texaschic101):

where are they ?

OpenStudy (anonymous):

is it ok that you guys help on checking my test right? i dont wanna break the rules

OpenStudy (texaschic101):

not really supposed to.....

OpenStudy (anonymous):

people do it anyways... youre ok

OpenStudy (anonymous):

ok thanks

OpenStudy (texaschic101):

we can probably check them and at least tell you if you got them right or wrong

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yeah thats what i want just dont say its A or its 3.6 k? thats the rule

OpenStudy (anonymous):

let me guess or try to solve it

OpenStudy (texaschic101):

thats a deal

OpenStudy (anonymous):

deal

OpenStudy (anonymous):

The rate table shows how many miles Emma can ride her bike at a constant rate. Complete the table.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):



OpenStudy (anonymous):

i say its 3,6,12,18

OpenStudy (texaschic101):

nothing but equivalent fractions 3/1 = x/2 equivalent fraction with 2 as denominator is 6/2.....3/1 = x/4...equivalent fraction with 4 as denominator is 12/4....just find the equivalent fractions

OpenStudy (anonymous):

yeah that's right because 1x3=3 2x3=6 4x3=12 and 6x3=18

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which equations could be used to solve the problem? Snow fell steadily, accumulating 14 in. in 4 h. At what rate was the snow falling? Choose exactly two answers that are correct. A. 4 h • r = 14 in. B. C. D. 14 inches • r = 4 h

OpenStudy (anonymous):

oops i have to type this one out

OpenStudy (anonymous):

A. 4h * r = 14in B. r = 4 h/14 in C. r = 14in/4h D. 14 in * r = 4h

OpenStudy (anonymous):

im not quite sur on this but i say its C

OpenStudy (anonymous):

you guys still there?

OpenStudy (anonymous):

=(